OpenStudy (anonymous):

Will give medal to the helper :) !! Simplify 2logx-log5=-2

OpenStudy (triciaal):

2 log x = log x^2 when you subtract logs you divide the numbers -2 can be expressed as the log of a certain number, use base 10 now what do you have?

OpenStudy (anonymous):

not sure

OpenStudy (triciaal):

use the info i gave you, take your time

OpenStudy (triciaal):

quick question was the 2 outside parentheses or just like you have it with the x only?

OpenStudy (anonymous):

just like i have it

OpenStudy (triciaal):

ok

OpenStudy (triciaal):

i have log (x^2/5) = -2 if Log (base 10) y = -2 what is Y?

OpenStudy (triciaal):

y would be 10^-2 = 1/100 do you agree?

OpenStudy (anonymous):

yes

OpenStudy (triciaal):

so from original log (x^2/5) = log (1/100) so x^2/5 = 1/100 solve for x

OpenStudy (anonymous):

how do i do it

OpenStudy (triciaal):

i really think it should be (x/5)^2 so that x^2/25 = 1/100 then x^2 = 25/100= 1/4 and x ==/- 1/2

OpenStudy (anonymous):

it is supposed to be .223

OpenStudy (triciaal):

well it is as written x^2/5 = 1/100 solve for x leave the variable on one side

OpenStudy (triciaal):

x^2 = 5/100 =0.05 x=rt 0,05 ==/-0.223

OpenStudy (anonymous):

k

OpenStudy (triciaal):

any ?

OpenStudy (triciaal):

typo == should be =+